OpenStudy (caroline1218):

Which equation of a line is perpendicular to the given line and passes through (-2, 3) y - 3 = 8/3(x +2)

OpenStudy (caroline1218):

The above equation is y = 8/3x + 25/3. Since it's slope is 8/3, the slope for the perpendicular line will be -3/8.

OpenStudy (caroline1218):

Is the correct answer y - 3 = -3/8(x + 2) OR y -2 = -3/8(x + 3)

OpenStudy (caroline1218):

@imqwerty @jhonyy9 @mathmale @nincompoop

OpenStudy (safa102):

i'm alittle confused on this!

jhonyy9 (jhonyy9):

y-3= 8/3(x+2) y= 8/3(x+2) +3 y = (8x+16 +9)/3 y = (8x +25)/3 y = 8x/3 +25/3